# references.bib

### v2

@@ B.3 Visualization of Sparse Semantic Features
Fig. 9 visualizes the sparse semantic features obtained from the dense encoder by projecting the 2D feature map to 3D form using the reference coordinates. The sparse semantic features are aligned to the input point cloud using weighted spatial interpolation, as detailed in §3.2. Although the 2D feature map output by the dense encoder is in low resolution (32$\times$18 for DINOv2 [dino] backbone), we still observe clear and distinguishable continuous feature variations on the aligned features, where the targets at the current step can be easily identified from the entire scene. Such a characteristic ensures precise feature fusion in the spatial domain. Additionally, we find that the features from the visual foundation model DINOv2 [dino] change significantly as the task progresses, enabling the model to clearly understand the global state at the current time.
@@ C.1 Setup
The teleoperated demonstrations are collected using AirExo [airexo_v1], while in-the-wild demonstrations are collected and adapted using our proposed AirExo-2 system. For each task, unless further specified, we collect 50 teleoperated demonstrations for RISE-2 policy evaluation and 50 in-the-wild demonstrations for AirExo-2 system evaluations.
@@
We compare RISE-2 against several representative policies based on 2D images and 3D point clouds, including: (1) ACT [act], which employs transformers to map image observations and proprioception to robot action chunks; (2) Diffusion Policy [dp], which formulates action prediction as a diffusion denoising process [ddpm, ddim] conditioned on the image observations; (3) CAGE [cage], an extension of Diffusion Policy that incorporates visual foundation models [dino], a causal observation perceiver [perceiver], and an attention-based diffusion action head for improved generalization; (4) RISE [rise], a 3D imitation policy that leverages a sparse 3D encoder for efficient point cloud perception; and (5) $\pi_{0}$ [pi0], a vision-language-action flow model for general robot control.
@@
All policies except $\pi_{0}$ are deployed on a workstation with an NVIDIA RTX 2060 SUPER GPU, while $\pi_{0}$ is deployed remotely on a server with an NVIDIA A100 GPU for GPU memory considerations. Following the procedure outlined in [umi, cage], we adopt a consistent evaluation method for each policy to minimize performance variation and ensure reproducibility. Specifically, we generate uniformly distributed test positions randomly before each task evaluation. The workspace is set up identically across different policies and test environments, and success rates are recorded for each test case. Each policy is evaluated over 20 consecutive trials per task (except for the Serve Steak task, where we evaluate each policy for 10 consecutive trials), and the success rates are computed accordingly.
@@ C.2 Action Accuracy Analysis Details
To evaluate the action accuracy of the data collection systems, we designed a special evaluation board with three tracks, as illustrated in Fig. 11. Each track has fixed holes spaced 2 cm apart. We created custom connectors for both the AirExo-2 and UMI [umi] that fit into these fixed holes, allowing us to collect position data. By sequentially placing the connectors into each fixed hole along the track, we can calculate the relative movement distance between two adjacent fixed holes and compare it with the true value (20 mm) to calculate the error.
@@ C.3 Case Study: Are In-Hand Cameras Sufficient for Manipulations?
We conducted a case study to investigate whether in-hand cameras are sufficient for many manipulation tasks. We select the Collect Toys task as an example and utilize CAGE [cage] as the policy for this case study.
Figure 12: Case Study Results. (Top) In-hand cameras often yield low-quality depth during object interaction, limiting their use for 3D point cloud-based policies. (Bottom) This case study uses CAGE with relative actions, as in the original implementation [cage]; subsequent experiments use CAGE with absolute actions.
In-hand cameras alone are often insufficient for manipulation tasks and may pose additional obstacles on policy learning. As shown in Fig. 12 (bottom), neither global nor in-hand cameras alone provide adequate observations for 2D image-based policies to achieve strong performance. Recent work [rise] has demonstrated that using only a global camera enables a 3D imitation policy to outperform 2D multi-view image-based policies, highlighting the importance of 3D information for scene understanding. However, as illustrated in Fig. 12 (top), in-hand cameras may produce incomplete depth information when the robotic arm approaches an object, making them unsuitable for 3D point-cloud-based policies. Consequently, relying solely on in-hand cameras can degrade the performance of the policies, particularly for 3D policies that rely on complete and accurate depth information to achieve superior performance.
@@ C.4 Additional Experiment Results
We also evaluate the performance of the RISE policy [rise] using pseudo-robot demonstrations. Specifically, we employ AirExo-2 to collect an equivalent number of in-the-wild demonstrations, which are then transformed into pseudo-robot data and used to train the RISE policy. The trained policy is deployed on a real robot without any fine-tuning. As shown in Fig. 13 and Tab. 7, RISE achieves good zero-shot performance, though with a slightly larger drop compared to RISE-2. These findings highlight the critical role of a generalizable policy in successfully transferring manipulation skills from in-the-wild demonstrations to real robots, under the circumstances of no teleoperation data.
@@
AirExo-2 RISE [rise] 60.625% -13.750%
[…]
Table 7: Evaluation Results of Different Policies in Learning from Pseudo-Robot Demonstrations. RISE-2 shows a smaller performance drop than RISE [rise] when trained with adapted pseudo-robot demonstrations.
@@
We conduct additional experiments on action representations for the Collect Toys task, comparing relative and absolute action representations. The results in Tab. 8 show that while relative action representation sometimes yields better results, absolute action representation provides more stable performance, particularly in terms of generalization. Therefore, we use absolute action representations for Diffusion Policy [dp] and CAGE [cage] throughout our experiments, except for the case study in Appendix C.3. For $\pi_{0}$ [pi0], we still use relative action representation for consistency with its released pre-trained models.
@@
Diffusion Policy [dp] relative 37.5% 20.0% 5.0%

### v3

@@ B.3 Visualization of Sparse Semantic Features
[…]
@@ C.1 Setup
The teleoperated demonstrations are collected using AirExo [29], while in-the-wild demonstrations are collected and adapted using our proposed AirExo-2 system. For each task, unless further specified, we collect 50 teleoperated demonstrations for RISE-2 policy evaluation and 50 in-the-wild demonstrations for AirExo-2 system evaluations.
@@
We compare RISE-2 against several representative policies based on 2D images and 3D point clouds, including: (1) ACT [123], which employs transformers to map image observations and proprioception to robot action chunks; (2) Diffusion Policy [18], which formulates action prediction as a diffusion denoising process [37, 92] conditioned on the image observations; (3) CAGE [111], an extension of Diffusion Policy that incorporates visual foundation models [75], a causal observation perceiver [44], and an attention-based diffusion action head for improved generalization; (4) RISE [102], a 3D imitation policy that leverages a sparse 3D encoder for efficient point cloud perception; and (5) $\pi_{0}$ [7], a vision-language-action flow model for general robot control.
@@
All policies except $\pi_{0}$ are deployed on a workstation with an NVIDIA RTX 2060 SUPER GPU, while $\pi_{0}$ is deployed remotely on a server with an NVIDIA A100 GPU for GPU memory considerations. Following the procedure outlined in [19, 111], we adopt a consistent evaluation method for each policy to minimize performance variation and ensure reproducibility. Specifically, we generate uniformly distributed test positions randomly before each task evaluation. The workspace is set up identically across different policies and test environments, and success rates are recorded for each test case. Each policy is evaluated over 20 consecutive trials per task (except for the Serve Steak task, where we evaluate each policy for 10 consecutive trials), and the success rates are computed accordingly.
@@ C.2 Action Accuracy Analysis Details
To evaluate the action accuracy of the data collection systems, we designed a special evaluation board with three tracks, as illustrated in Fig. 11. Each track has fixed holes spaced 2 cm apart. We created custom connectors for both the AirExo-2 and UMI [19] that fit into these fixed holes, allowing us to collect position data. By sequentially placing the connectors into each fixed hole along the track, we can calculate the relative movement distance between two adjacent fixed holes and compare it with the true value (20 mm) to calculate the error.
@@ C.3 Case Study: Are In-Hand Cameras Sufficient for Manipulations?
We conducted a case study to investigate whether in-hand cameras are sufficient for many manipulation tasks. We select the Collect Toys task as an example and utilize CAGE [111] as the policy for this case study.
Figure 12: Case Study Results. (Top) In-hand cameras often yield low-quality depth during object interaction, limiting their use for 3D point cloud-based policies. (Bottom) This case study uses CAGE with relative actions, as in the original implementation [111]; subsequent experiments use CAGE with absolute actions.
In-hand cameras alone are often insufficient for manipulation tasks and may pose additional obstacles on policy learning. As shown in Fig. 12 (bottom), neither global nor in-hand cameras alone provide adequate observations for 2D image-based policies to achieve strong performance. Recent work [102] has demonstrated that using only a global camera enables a 3D imitation policy to outperform 2D multi-view image-based policies, highlighting the importance of 3D information for scene understanding. However, as illustrated in Fig. 12 (top), in-hand cameras may produce incomplete depth information when the robotic arm approaches an object, making them unsuitable for 3D point-cloud-based policies. Consequently, relying solely on in-hand cameras can degrade the performance of the policies, particularly for 3D policies that rely on complete and accurate depth information to achieve superior performance.
@@ C.4 Additional Experiment Results
We also evaluate the performance of the RISE policy [102] using pseudo-robot demonstrations. Specifically, we employ AirExo-2 to collect an equivalent number of in-the-wild demonstrations, which are then transformed into pseudo-robot data and used to train the RISE policy. The trained policy is deployed on a real robot without any fine-tuning. As shown in Fig. 13 and Tab. 7, RISE achieves good zero-shot performance, though with a slightly larger drop compared to RISE-2. These findings highlight the critical role of a generalizable policy in successfully transferring manipulation skills from in-the-wild demonstrations to real robots, under the circumstances of no teleoperation data.
@@
AirExo-2 RISE [102] 60.625% -13.750%
RISE-2 85.000%  $^{\text{ }\underline{-8.750\%}}$
Table 7: Evaluation Results of Different Policies in Learning from Pseudo-Robot Demonstrations. RISE-2 shows a smaller performance drop than RISE [102] when trained with adapted pseudo-robot demonstrations.
@@
We conduct additional experiments on action representations for the Collect Toys task, comparing relative and absolute action representations. The results in Tab. 8 show that while relative action representation sometimes yields better results, absolute action representation provides more stable performance, particularly in terms of generalization. Therefore, we use absolute action representations for Diffusion Policy [18] and CAGE [111] throughout our experiments, except for the case study in Appendix C.3. For $\pi_{0}$ [7], we still use relative action representation for consistency with its released pre-trained models.
@@
Diffusion Policy [18] relative 37.5% 20.0% 5.0%
@@
C.5 Failure Analysis
Imprecise Action.
During the policy rollouts, we observe that most failures stem from imprecise predicted actions. For example, in the Open Lid task, the lid protrudes only about 1.5 cm from the coffee machine body, requiring highly accurate end-effector control from the visuomotor policies. At the same time, the left gripper must stabilize the machine to prevent it from shifting on the table while the right arm applies force to lift the lid. Baseline failures typically arise from inaccuracy in either arm — where the left arm fails to stabilize the coffee machine or the right arm fails to open the lid. In contrast, our RISE-2 policy produces more precise actions, resulting in fewer failures, even when trained with adapted in-the-wild data.
Risk of Visual Overfitting to Robot States.
We also observe that in bimanual manipulation tasks, since both arms occupy a large portion of the observation image, there is an increased risk of visuomotor policies overfitting to robot states. This occurs even without proprioceptive inputs, as the policy may implicitly infer actions from the visual configuration of the arms. The issue is further compounded by limited training data: with only 50 demonstrations, most baselines fail to learn sufficiently accurate actions, leading to low phase success rates. Such overfitting also reduces the likelihood of performing recovery behaviors after phase failures, which becomes especially evident in long-horizon and complex tasks such as Serve Steak, ultimately causing frequent task failures. In contrast, RISE-2 produces more precise actions, resulting in higher success rates. Nevertheless, all policies still exhibit varying degrees of overfitting to robot visual state and struggle to recover from failures in the absence of recovery demonstrations, highlighting an important direction for developing more generalizable robotic visuomotor policies.
Appendix D Author Contributions
The author contributions for this paper are as follows. For each item, authors are listed in order of contribution, and underlined names indicate equal contributions.
  • Project Lead: Hongjie Fang.
  • System Design and Implementation.
    • AirExo-2 Hardware: Yiming Wang, Hongjie Fang, Hao-Shu Fang.
    • AirExo-2 Visual Adaptors: Jingjing Chen, Hongjie Fang.
    • RISE-2 Policy: Chenxi Wang, Hongjie Fang, Jingjing Chen, Shangning Xia.
  • Experiments:
    • Design: Hongjie Fang, Chenxi Wang, Jingjing Chen, Yiming Wang, Hao-Shu Fang.
    • Platform Setup: Hongjie Fang, Yiming Wang.
    • Data Collection: Hongjie Fang, Chenxi Wang, Jingjing Chen.
    • Data Processing: Jingjing Chen, Zihao He, Xiyan Yi, Yunhan Guo.
    • Baselines: Hongjie Fang, Shangning Xia, Chenxi Wang, Jingjing Chen.
    • Policy Training: Hongjie Fang, Chenxi Wang, Shangning Xia, Jingjing Chen, Xinyu Zhan, Lixin Yang.
    • Policy Evaluation: Hongjie Fang, Chenxi Wang, Jingjing Chen.
    • User Study: Hongjie Fang, Chenxi Wang, Jingjing Chen, Xiyan Yi.
    • Accuracy Analysis: Hongjie Fang, Jun Lv, Jingjing Chen, Yiming Wang.
    • Scalability Analysis: Hongjie Fang, Jingjing Chen, Chenxi Wang, Hao-Shu Fang.
  • Writing: Hongjie Fang, Chenxi Wang, Jingjing Chen, Yiming Wang, Hao-Shu Fang.
  • Project Advisor: Hao-Shu Fang, Cewu Lu, Weiming Wang.